UITF Meeting - March 4, 2019
UITF schedule 3/4/2019
- Run keV beam, compare gun performance tee-shaped electrode versus modified mushroom (now through 2019)
- Continue testing the tee-shaped electrode at 200kV (the highest voltage polarized source in the world!) We have been measuring bunch length now using choppers and MS Initial bunch length measurements changing laser phase and then chopper phase
- Larry Farrish and Mark Wissman worked on Buncher, seems ready. Now need Rama and/or Clyde to setup EPICS controls and calibrate.
- Test the modified mushroom that Bubba just finished polishing CEBAF small shield polished electrode - waiting on internal electrode parts due end of March
- Then in august, we install this modified mushroom electrode at CEBAF
- Then we have to modify the big mushroom electrode we pull from CEBAF and install it at UITF – this will take weeks, probably can’t wait that long, so guessing we re-install the tee? More bakeouts!
- The tasks listed above represent Gabriel’s thesis work, how about Josh beam studies?, should not be a problem
- Beam to Brock cavity polarimeter, needs to happen, Brock wants to be here for it (that means venting the spectrometer dump, installing cavity and re-baking it, sigh)
- Float beam through QCM, test the BCM electronics (RF OFF, valves OPEN)
- Fix the north wall – SEG blocks, roof tile, helium chimney vent tile, re-work the electric conduit
- Facilities claims this work to be complete mid-April, some of it will undoubtedly interrupt keV operations
- John Riesbeck will be coordinating this work. Electrical work was supposed to have been started last week.
- Final RadCon shielding documentation
- Need to finish the MeV beamline, and provide all the controls. We have our stuff under control, but what about Engineering, can they get this done by end of March? Most of this work interrupts keV beam operations. We have now a 'routine' of leaving the cave OPEN in morning, and locking it up in afternoons
- Vacuum work, pump down, leak check, stack lead
- Harps, viewers, current monitoring, BPMs, valves, ions pumps, yao cavities and electronics, Update from I&C?
- Magnets, third rack: order haimsons (long lead time!, I’m on it), pull magnets from LERF. Status of third magnet rack?
- as-built song sheets complete, UED updated, MPS fully functional, channel access finalized
- Approval to make MeV beam
- Roof tile OSP (writing it now)
- MeV OSP
- 3rd review, conduct of operations review, UITF AOD (April)
- ARR (mid-May? How unrealistic is that?)
- QCM commissioning, HDIce tests
- Finish the HDIce beamline, upstream and downstream, raster magnet
- Reza, Yan, and CIS, define the metrics for good beam and good QCM
- Fight for LHe!! Week-long run periods, staffing?